Please help me solve this problem:
If you earn $3,000 over 10 years on an investment that pays 5% compounded annually, what was the principle amount you started with?

Found 2 solutions by ewatrrr, mananth:
Answer by ewatrrr(24785)


Hi
Earning $3,000 over 10 years on an investment that pays 5% compounded 'annually'
Note: A= P(1+r/n) ^nt OR in this case: P + I = P(1+r) ^t and I = P[(1+r)^t - 1]
$3000 = P[(1.05)^10 - 1]
$3000/.6289 = P
$4770.23 = p
CHECKING our Answer***
$7770.23 = $4770.23(1.05)^10

Answer by mananth(16946)

A=p(1+r/n)^nt
p+3000=p(1+0.05/10)^100
p+3000+1.65p
0.65p=3000
p=3000/0.65
P= $4,675